sonic boom You hear the sonic boom of a high-speed jet plane exactly 4.30 s after it passes directly overhead in level flight. At the time you hear the boom, you see the plane at an angle of 30.0 o above the horizon. Assume that the speed of sound at the altitude of the plane is 325 m/s. How fast is the plane traveling? What is the altitude of the plane?
Answer #1

Let the plane be at an altitude 'h' meters with a speed of 'v' m/s

The sound travelled directly downwards towards the person, for sound to reach the person,

h = 325*4.3 (distance = speed*time)

h = 1397.5 meters

Horizontal distance covered by the plane in this time interval,

d = v*4.3

Given, =30o

and , tan() = d/h

tan(30) = 4.3*v/1397.5

v = 187.63 m/s